As a member of Engineers Without Borders, you are working in a community that has contaminated drinking water caused by slow just answer part 3
Answer #1

Since the coefficient of determination (r2) is more for the power model (1) than the exponential model (.905) we say that the power model fits the data the best.

The power model is given as y^= a* xb = 280.184 * x.12

For bacteria concentration to reach 475/1000mL,the time taken by power model is given as 475 = 280.184 * x.12

Then x= 81.36 = 81 hours (app)